### v2

# MiLoRA: Harnessing Minor Singular Components for Parameter-Efficient LLM Finetuning

Abstract
Efficient finetuning of large language models (LLMs) aims to adapt the LLMs with reduced computational and memory cost. Previous LoRA-based approaches initialize the low-rank matrices with Gaussian distribution and zero values while keeping the original weight matrices frozen. However, the trainable model parameters optimized in an unguided subspace might interfere with the well-learned subspace of the pretrained weight matrices. In this paper, we propose MiLoRA, a simple yet effective LLM finetuning approach that only updates the minor singular components of the weight matrix while keeping the principal singular components frozen. It is observed that the minor matrix corresponds to the noisy or long-tail information, while the principal matrix contains important knowledge. The MiLoRA initializes the low-rank matrices within a subspace that is orthogonal to the principal matrix, thus the pretrained knowledge is expected to be well preserved. During finetuning, MiLoRA makes the most use of the less-optimized subspace for learning the labeled dataset. Extensive experiments on commonsense reasoning, math reasoning, instruction following and visual instruction following benchmarks present the superior performance of our method.
MiLoRA: Harnessing Minor Singular Components for Parameter-Efficient LLM Finetuning
Hanqing Wang1,  Yixia Li2,  Shuo Wang3,  Guanhua Chen2,  Yun Chen1 1Shanghai University of Finance and Economics 2Southern University of Science and Technology   3Tsinghua University
1 Introduction
Large language models (Brown et al., 2020; Ouyang et al., 2022; Touvron et al., 2023; Jiang et al., 2023, LLMs) have demonstrated superior performance on various tasks (Zheng et al., 2023), such as math reasoning (Wang et al., 2024a) and question answering (Ivison et al., 2023). These models are pretrained with the next token prediction task (Brown et al., 2020) on large web-scale data, then finetuned with instruction data as well as human preference data (Ouyang et al., 2022; Yu et al., 2023; Cui et al., 2023) for different downstream tasks. Fully finetuning is commonly employed to unlock the complete potential of LLMs, however, optimizing all model parameters necessitates substantial and restrictive computational resources (Touvron et al., 2023; Jiang et al., 2023), which hampers the utilization of LLMs across diverse scenarios.
Parameter-efficient finetuning (Hu et al., 2021, 2023; Liu et al., 2022; Zhang et al., 2024; Liu et al., 2024b, PEFT) aims at reducing the computational and GPU memory cost for finetuning of pretrained models. Low-rank adaptation (Hu et al., 2021, LoRA) is one of the most widely used PEFT methods for LLM finetuning. It assumes the change of linear model weights to be low-rank (Li et al., 2018; Aghajanyan et al., 2021). For each selected weight matrix, it only updates two low-rank matrices while keeping the pretrained weight frozen. During inference, the low-rank matrices are merged into the pretrained linear weights, thus no additional computational or memory cost is introduced. Recently, researchers have explored different LoRA-based variants for efficient LLM finetuning (Zhang et al., 2023; Pan et al., 2024; Kopiczko et al., 2024; Liu et al., 2024b; Meng et al., 2024). However, most existing LoRA-based methods randomly initialize the low-rank matrices and optimize the trainable parameters in an unguided subspace. We suspect this strategy may override important pretrained features, thus degrading the performance of low-rank adaptation methods (Dou et al., 2024).
In this paper, we propose Minor singular component based Low Rank Adaptation (MiLoRA) for efficient LLM finetuning. MiLoRA has a similar framework as LoRA but employs a different initialization schedule. Specifically, a weight matrix ${\bm{W}}$ is decomposed with the singular value decomposition (SVD) algorithm. Based on the magnitude of the singular values, we divide ${\bm{W}}$ into two components: the principal matrix ${\bm{W}}_{p}$ corresponding to large singular values and the minor matrix ${\bm{W}}_{m}$ corresponding to small singular values. We argue that the principal matrix captures the essence of the pretrained knowledge, whereas the minor matrix is suboptimal with noisy or long-tail information. It is supported by previous works (Hajimolahoseini et al., 2021; Sharma et al., 2024) that principal low-rank approximation can achieve comparable or even better performance to full parameters.
Motivated by these observations, we freeze the principal matrix ${\bm{W}}_{p}$ and adapt the minor singular components during finetuning. The low-rank matrices ${\bm{A}}$ and ${\bm{B}}$ in LoRA framework are initialized with the minor matrix ${\bm{W}}_{m}$. Since the trainable low-rank matrices are initialized in a subspace orthogonal to the principal matrix, MiLoRA is expected to effectively learn from finetuning tasks while better preserving and utilizing the pretrained knowledge. To maintain the capability of the pretrained model at the start of finetuning, vanilla LoRA explicitly initializes ${\bm{B}}$ with zeros. In contrast, MiLoRA naturally satisfies this requirement, as the pretrained weight matrix ${\bm{W}}$ equals the frozen principal part ${\bm{W}}_{p}$ plus the low-rank part ${\bm{W}}_{m}={\bm{B}}{\bm{A}}$. We conduct extensive experiments on commonsense reasoning, math reasoning, instruction following and visual instruction following benchmarks. The experimental results show that MiLoRA consistently outperforms LoRA without sacrificing training or inference efficiency, such as commonsense reasoning (+1.6/+1.1 on LLaMA2-7B/LLaMA3-8B), math reasoning (+2.0 on LLaMA2-7B), instruction following (+1.4 on LLaMA2-7B) and visual instruction following (+1.4 on LLaVA1.5-7B). [^1] Our code and model will be made public.
2 Preliminaries
Singular Value Decomposition
Given a matrix ${\bm{W}}\in\mathbb{R}^{m\times n}$, its singular value decomposition is denoted as ${\bm{W}}={\bm{U}}\bm{\Sigma}{\bm{V}}^{\top}$, where ${\bm{U}}=\left[u_{1},u_{2},\cdots,u_{m}\right]\in\mathbb{R}^{m\times m}$, ${\bm{V}}=\left[v_{1},v_{2},\cdots,v_{n}\right]\in\mathbb{R}^{n\times n}$. The columns of ${\bm{U}}$ are the left singular vectors, and the columns of ${\bm{V}}$ are the right singular vectors. The $\bm{\Sigma}\in\mathbb{R}^{m\times n}$ is a diagonal matrix containing the singular values of ${\bm{W}}$ in descending order.
Without loss of generality, we suppose $m\leq n$ to simplify the notation. The SVD of ${\bm{W}}$ can be reformulated as in Equation 1.
${\bm{W}}={\bm{U}}\bm{\Sigma}{\bm{V}}^{\top}=\sum_{i=1}^{m}\sigma_{i}u_{i}v_{i}^{\top},$ (1)
where $u_{i}$ and $v_{i}$ are the $i^{\mathrm{t}h}$ column of ${\bm{U}}$ and ${\bm{V}}$, respectively.
Low-Rank Adaptation
The low-rank adaptation method (Hu et al., 2021, LoRA) assumes the updates of linear weight ${\bm{W}}\in\mathbb{R}^{m\times n}$ to be low-rank, thus models the changes with two trainable low-rank matrices ${\bm{A}}\in\mathbb{R}^{r\times n}$ and ${\bm{B}}\in\mathbb{R}^{m\times r}$. The weight matrix can be decomposed as
${\bm{W}}={\bm{W}}^{(0)}+\Delta{\bm{W}}={\bm{W}}^{(0)}+\frac{\alpha}{r}{\bm{B}}{\bm{A}},$ (2)
where ${\bm{W}}^{(0)}$ and $\Delta{\bm{W}}$ refer to the pretrained weight and weight change, respectively. The $\alpha$ and $r$ are hyperparameters of scaling-factor and LoRA rank ($r\ll\min(m,n)$). During finetuning, the pretrained matrix ${\bm{W}}^{(0)}$ is kept frozen. It significantly diminishes the number of trainable parameters as both ${\bm{A}}$ and ${\bm{B}}$ matrices are low-rank. The ${\bm{B}}$ matrix is initialized with zero while ${\bm{A}}$ matrix adopts a random Gaussian distribution with zero mean value.
This initialization strategy ensures the $\Delta{\bm{W}}=0$ at the beginning of training. The LoRA method only modifies the linear matrices in the Transformer model. The low-rank matrices can be easily merged into the pretrained linear matrix to get updated for inference, which does not require additional computing and GPU memory compared with full finetuning. However, the vanilla LoRA method fails to select the optimal subspace for updating the model parameters, as the low-rank matrices ${\bm{A}}$ and ${\bm{B}}$ are randomly initialized. This might potentially detract from the pretrained knowledge encoded in the pretrained weight matrix.
Figure 1: (a) MiLoRA method splits the pretrained linear weight matrix into two parts, the minor singular component (${\bm{U}}_{m},{\bm{V}}_{m},{\bm{\Sigma}}_{m}$) is used to initialize the low-rank matrices ${\bm{A}}_{m}$ and ${\bm{B}}_{m}$ in the LoRA framework. (b) During finetuning, only the low-rank matrices ${\bm{A}}_{m}$ and ${\bm{B}}_{m}$ are updated while the principal matrix ${\bm{W}}_{p}$ is frozen, which is similar to the practice of LoRA method.
3 Methodology
The proposed MiLoRA is a simple yet effective PEFT approach that selects the minor singular components of the weight matrices for optimization. As observed in LASER (Sharma et al., 2024), the minor singular components of weight matrices contain noisy or long-tail information, while the principal singular components are responsible for important features across tasks. Therefore, MiLoRA is designed to effectively learn from the finetuning dataset by adapting minor singular components, while maintaining the knowledge encoded in the pretrained model. On the contrary, vanilla LoRA fails to constrain the optimization subspace as its low-rank matrices are randomly initialized. Specifically, suppose $m\leq n$, MiLoRA splits each linear weight matrix into two matrices according to the corresponding singular values: principal matrix ${\bm{W}}_{p}$ and minor matrix ${\bm{W}}_{m}$ (see Figure 1):
${\bm{W}}$ $={\bm{W}}_{p}+{\bm{W}}_{m}$ (3)
$=\sum_{i=1}^{m-r}\sigma_{i}u_{i}v_{i}^{\top}+\sum_{i=m-r+1}^{m}\sigma_{i}u_{i}v_{i}^{\top},$
where the singular values are in descending order ($\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{m}\geq 0$), and the hyperparameter $r$ is the number of minor singular values in the ${\bm{W}}_{m}$ matrix.
The decomposition can be also reformulated in a matrix form. The ${\bm{U}}$ matrix in the SVD of ${\bm{W}}$ can be reformulated as ${\bm{U}}=\left[{\bm{U}}_{p},{\bm{U}}_{m}\right]$, where ${\bm{U}}_{p}=\left[u_{1},u_{2},\cdots,u_{m-r}\right]$ and ${\bm{U}}_{m}=\left[u_{m-r+1},u_{m-r+2},\cdots,u_{m}\right]$ are left singular vectors corresponding to principal and minor singular values, respectively. The ${\bm{V}}$ and ${\bm{\Sigma}}$ matrices can be reformulated similarly. The SVD of ${\bm{W}}$ can be expressed as:
${\bm{W}}$ $={\bm{U}}{\bm{\Sigma}}{\bm{V}}^{\top}={\bm{U}}_{p}{\bm{\Sigma}}_{p}{\bm{V}}_{p}^{\top}+{\bm{U}}_{m}{\bm{\Sigma}}_{m}{\bm{V}}_{m}^{\top}$ (4)
$={\bm{W}}_{p}+{\bm{W}}_{m}$
During the finetuning process, instead of freezing the entire pretrained weight matrix, we just keep the principal singular components ${\bm{W}}_{p}$ fixed to preserve the pretrained knowledge. The minor matrix ${\bm{W}}_{m}$ is used to initialize low-rank matrices ${\bm{A}}_{m}$ and ${\bm{B}}_{m}$ (see Figure 1):
${\bm{W}}_{m}$ $={\bm{U}}_{m}{\bm{\Sigma}}_{m}{\bm{V}}_{m}^{\top}$ (5)
$=({\bm{U}}_{m}\sqrt{{\bm{\Sigma}}_{m}})(\sqrt{{\bm{\Sigma}}_{m}}{\bm{V}}_{m}^{\top})={\bm{B}}_{m}{\bm{A}}_{m},$
This strategy has two benefits: 1) It encourages the model to learn in the less-optimized subspace spanned by the minor singular vectors, thus reducing the interference with the well-learned pretrained knowledge encoded in the principal singular components. 2) Unlike vanilla LoRA which requires tuning the scaling factor $\alpha$ and the initialization hyperparameters, ours does not require any hyperparameter except the rank $r$.
To make our method simple and training-efficient, we do not incorporate the regularization loss which imposes strict constraints on the orthogonality between ${\bm{W}}_{m}$ and ${\bm{W}}_{p}$. Our experiment results show that our scheme works quite well in balancing learning from the finetuning dataset and preserving pretrained knowledge.
4 Experiments
Model PEFT BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c OBQA Avg.
ChatGPT -- 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
LLaMA2-7B LoRA 69.8 79.9 79.5 83.6 82.6 79.8 64.7 81.0 77.6
PiSSA 67.6 78.1 78.4 76.6 78.0 75.8 60.2 75.6 73.8
MiLoRA 67.6 83.8 80.1 88.2 82.0 82.8 68.8 80.6 79.2
LLaMA3-8B LoRA 70.8 85.2 79.9 91.7 84.3 84.2 71.2 79.0 80.8
PiSSA 67.1 81.1 77.2 83.6 78.9 77.7 63.2 74.6 75.4
MiLoRA 68.8 86.7 77.2 92.9 85.6 86.8 75.5 81.8 81.9
Table 1: Commonsense reasoning evaluation results for LLaMA2-7B and LLaMA3-8B on eight tasks. Results are cited from Liu et al. (2024b). Bold numbers indicate the highest performance scores for each dataset across the different PEFT methods for the corresponding model.
Method GSM8K MATH Avg.
Full FT 66.5 19.8 43.2
LoRA 60.6 16.9 38.7
PiSSA 58.2 15.8 37.0
MiLoRA 63.5 17.8 40.7
Table 2: Math reasoning evaluation results for LLaMA2-7B on GSM8K and MATH. Results are cited from Yu et al. (2023).
To evaluate MiLoRA against other PEFTs, we conduct experiments on both NLP and multimodal tasks covering 18 datasets. We benchmark against LoRA (Hu et al., 2021) and PiSSA (Meng et al., 2024). All of our experiments are conducted on 8 NVIDIA L40 GPUs unless otherwise specified.
  • LoRA (Hu et al., 2021) reparameterizes the weight update $\Delta{\bm{W}}$ with two trainable low-rank matrices ${\bm{A}}$ and ${\bm{B}}$, while freezing the pretrained weight ${\bm{W}}$. They use a random Gaussian initialization for A and zero for B, so $\Delta{\bm{W}}={\bm{B}}{\bm{A}}$ is zero at the beginning of training.
  • PiSSA (Meng et al., 2024) shares the same framework as LoRA, but employs a more sophisticated initialization approach. They initialize ${\bm{A}}$ and ${\bm{B}}$ with principal singular values and singular vectors of the pre-trained weight ${\bm{W}}$. Given that the principal components capture the essence of a matrix, PiSSA is expected to better approximate full finetuning by changing the essential parts while freezing the “noisy” parts.
4.1 Experiments on Large Language Model
Models and Datasets
We evaluate and compare MiLoRA with baselines on three different types of LLM downstream tasks.
1. Commonsense reasoning: We finetune LLaMA2-7B (Touvron et al., 2023) and LLaMA3-7B (AI@Meta, 2024) on Commonsense170K (Hu et al., 2023). Eight commonsense reasoning datasets are used for evaluation, including BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-e, ARC-c (Clark et al., 2018), and OBQA (Mihaylov et al., 2018). The task is formulated as a multiple-choice problem. We report accuracy (%) for all datasets on the best checkpoint chosen by the validation set loss.
2. Math reasoning: We finetune LLaMA2-7B (Touvron et al., 2023) on the MetaMathQA dataset (Yu et al., 2023), which contains 395K samples augmented from the training set of GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021).
We use test sets of GSM8K and MATH for evaluation and report results on the last checkpoint. For each test set, we report the exact match ratio against the ground truth.
3. Instruction following: We follow the experiment setup in Wu et al. (2024a) and Wu et al. (2024b) to finetune LLaMA2-7B with Ultrafeedback (Cui et al., 2023). Following previous works, we use Alpaca-Eval v1.0 (Li et al., 2023a) for evaluation and report win-rate against text-davinci-003 using GPT4-0613 as the annotator (OpenAI, 2024).
Implementation Details
We use the same hyperparameter configurations as Hu et al. (2023) without tuning for all methods. Details can be found in Appendix A.2. We denote this hyperparameter setup as LLM-Adapters. By default, we use a rank of 32. As math reasoning has a large training dataset, we set the rank to 64.
Method Win-rate
GPT-4-0613 91.2
GPT-3.5-turbo-0613 91.8
LLaMA2-Chat-7B 87.3
LLaMA2-7B + LoRA 85.9
LLaMA2-7B + PiSSA 85.6
LLaMA2-7B + MiLoRA 87.3
Table 3: Instruction tuning evaluation results for instruction-tuned LLaMA2 7B with Alpaca-Eval v1.0. Results are cited from Ivison et al. (2023).
Results
Method VQAv2 GQA VisWiz SQA VQAT POPE MMBench Avg.
Full FT 78.5 61.9 50.0 66.8 58.2 85.9 64.3 66.5
LoRA 79.1 62.9 47.8 68.4 58.2 86.4 66.1 66.9
PiSSA 77.5 60.6 39.2 67.3 54.3 87.0 61.4 63.9
MiLoRA 79.2 62.1 53.3 70.6 58.7 87.9 66.1 68.3
Table 4: Visual instruction tuning evaluation results for LLaVA1.5-7B on seven vision-language tasks. We directly use checkpoints from Liu et al. (2024a) to reproduce the results of baseline methods.
We report results in Table 1, Table 2 and 3 for commonsense reasoning, math reasoning and instruction following, respectively. For commonsense reasoning, we also include the ChatGPT baseline reported in Liu et al. (2024b) as a reference, which is obtained with gpt-3.5-turbo API using a zero-shot Chain of Thought (Wei et al., 2022). For math reasoning, we add the full finetuning results from Yu et al. (2023) as a reference. For instruction following, we include the results of GPT-4-0613, GPT-3.5-turbo-0613, and LLaMA2-Chat-7B from Ivison et al. (2023) as a reference.
As can be seen, MiLoRA consistently surpasses all baseline methods across different datasets and LLMs, indicating that MiLoRA is a highly effective finetuning method. Specifically, for commonsense reasoning, MiLoRA outperforms LoRA and PiSSA by an average of 1.6 and 5.4 accuracy points on LLaMA2-7B, and by 1.1 and 6.5 points on LLaMA3-8B. For math reasoning, MiLoRA improves over LoRA and PiSSA by 2.0 and 3.7 averaged EM points on LLaMA2-7B, respectively. However, it still underperforms full finetuning by an average of 2.5 EM score. This suggests that the PEFT methods still have room for improvement to fully match the performance of full finetuning. For instruction following, MiLoRA achieves the best results, surpassing LoRA and PiSSA by 1.4 and 2.7 in win-rate scores, respectively.
4.2 Experiments on Vision-Language Model
Models and Datasets
To further examine if MiLoRA can remain competitive on multimodal finetuning tasks, we further conduct visual instruction tuning tasks on LLaVA1.5-7B Liu et al. (2024a), which consists of an LLM, Vicuna-1.5-7B Peng et al. (2023) and a vision encoder, CLIP ViT-L/336px Radford et al. (2021). The visual instruction tuning is performed in a multitask setting, including VQA Hudson and Manning (2019); Marino et al. (2019); Schwenk et al. (2022), OCR Mishra et al. (2019); Sidorov et al. (2020), region-level VQA Krishna et al. (2017); Mao et al. (2016), visual conversation Liu et al. (2024a), and language conversation data. The finetuned models are evaluated on seven different multimodal benchmarks: VQA-v2 Goyal et al. (2017), GQA Hudson and Manning (2019), VizWiz Gurari et al. (2018), SQA Lu et al. (2024), VQAT Singh et al. (2019), POPE Li et al. (2023b), and MMBench Liu et al. (2024c).
Implementation Details
We follow the settings of Liu et al. (2024a) to construct the training data and prompt template. For the hyperparameter configuration, we follow the LoRA configuration provided by Liu et al. (2024a) without tuning for all PEFT methods for a fair comparison. Details can be found in Appendix A.2.
Results
Table 4 shows the experiment results. As can be seen, MiLoRA achieves the best results, outperforming LoRA and PiSSA by 1.4 and 4.4 average accuracy points, respectively. We also note that in this setup, fully fine-tuning is less effective than LoRA. As a result, methods like PiSSA, which aim to closely approximate fully fine-tuning, may lose their advantage. In contrast, MiLoRA improves performance by retaining more of the pre-trained knowledge, thereby circumventing this issue.
5 Understanding MiLoRA
In this section, we conduct experiments to further understand MiLoRA. By default, we use the finetuned LLaMA2-7B model on math reasoning from our main experiments for analyses.
5.1 Comparison with More PEFTs
To further demonstrate the effectiveness of MiLoRA, we compare MiLoRA with additional popular baselines, such as rsLoRA Kalajdzievski (2023), LoRA+ Hayou et al. (2024), DoRA Liu et al. (2024b), AdaLoRA Zhang et al. (2023), LoRA-GA Wang et al. (2024b). We follow the math and code experiment settings in Wang et al. (2024b). Specifically, we finetune LLaMA2-7B with MiLoRA on MetaMathQA 100K and Code-Feedback 100K for one epoch using rank $8$ and compare MiLoRA with baseline results reported in Wang et al. (2024b). Table 5 shows the results. We report the training cost of all methods relative to LoRA, which requires 3.58 hours of training on the math task using a single NVIDIA-L40 GPU. As can be seen, MiLoRA consistently outperforms all baselines, improving the best PEFT baseline LoRA-GA by 1.1 and 4.6 points on GSM8K and Human-eval, respectively. The training efficiency of MiLoRA is also very high. The time for SVD decomposition in MiLoRA is less than six minutes and negligible when compared with the total training time. Therefore, the training time required of MiLoRA is consistent with that of LoRA, while DoRA and AdaLoRA require additional training time overhead.
Method GSM8K Human-eval Training Cost
LoRA 42.1 14.8 $\times$ 1.00
PiSSA 44.5 16.0 $\times$ 1.00
rsLoRA 45.6 16.0 $\times$ 1.00
LoRA+ 52.1 18.2 $\times$ 1.00
DoRA 53.1 19.8 $\times$ 1.58
AdaLoRA 50.7 17.8 $\times$ 1.07
LoRA-GA 53.6 19.8 $\times$ 1.00
MiLoRA 54.7 24.4 $\times$ 1.00
Table 5: Comparison with more baselines on math and code tasks. Baseline GSM8K and Human-eval results are cited from Wang et al. (2024b).
5.2 Is Minor Singular Component Adaptation Important?
To investigate the influence of singular components of varying magnitudes on finetuning performance, we conduct experiments on math reasoning tasks using the LLaMA2-7B model. Specifically, we finetune LLaMA2-7B on the MetaMathQA 395K dataset for 1 epoch with a maximum sequence length of 512 to save computation. We initialize the low-rank matrices ${\bm{A}}$ and ${\bm{B}}$ with principal, minor, and randomly sampled singular components, and report the evaluation results in Table 6. As can be seen, initializing with the minor singular components achieves the best performance across both datasets. This underscores the importance of adapting the minor singular components during the finetuning process. Concurrent with our work, Meng et al. (2024) propose PiSSA, a similar approach to MiLoRA but adapts the principal singular components. We will provide a detailed comparison between MiLoRA and PiSSA in Section 5.5.
Method principal Random Minor
GSM8K 60.7 63.2 64.0
MATH 14.6 15.5 16.1
Table 6: Performance of MiLoRA when initializing with principal, random sampled, and minor singular components.
5.3 How does the Matrix Update $\Delta W$ Compare to $W$?
We investigate the relationship between $\Delta{\bm{W}}$ and ${\bm{W}}$ of different methods, where $\Delta{\bm{W}}$ is the difference between the finetuned and pre-trained weights. Following the analysis method in Hu et al. (2021), we project ${\bm{W}}$ and $\Delta{\bm{W}}$ onto the $r$-dimensional subspace of $\Delta{\bm{W}}$ by computing ${\bm{U}}^{T}{\bm{W}}{\bm{V}}$ and ${\bm{U}}^{T}\Delta{\bm{W}}{\bm{V}}$, with ${\bm{U}}$/${\bm{V}}$ being the left/right top $r$ singular-vector matrix of $\Delta{\bm{W}}$. As a comparison, we also compute ${\bm{U}}^{T}{\bm{W}}{\bm{V}}$ by replacing ${\bm{U}}$, ${\bm{V}}$ with the top $r$ singular vectors of ${\bm{W}}$ or a random matrix. We use the Frobenius norm as the measurement of magnitude. The results are shown in Table 7. Following Hu et al. (2021), we use the query weight in the middle layer (16th) of the model for analysis. We make two key observations. First, $\Delta{\bm{W}}$ of LoRA and MiLoRA have a similar correlation with ${\bm{W}}$, both slightly stronger than a random matrix, indicating that they amplify directions not emphasized in ${\bm{W}}$. In contrast, $\Delta{\bm{W}}$ of PiSSA shows a strong correlation with ${\bm{W}}$. This is attributed to PiSSA’s direct optimization of the principal singular components. Second, the amplification factor varies significantly across different methods. Compared to LoRA, we find the amplification factor of MiLoRA is significantly smaller. This indicates that MiLoRA has a reduced impact on the features which already present in ${\bm{W}}$. See Table 13 of Appendix B for analyzing the weight matrix of MLP down projection in the same layer, which demonstrates similar trends.
$\left\|{\bm{W}}\right\|_{F}=97.36$ ${\bm{W}}$ Random LoRA PiSSA MiLoRA
$\left\|{\bm{U}}^{\top}{\bm{W}}{\bm{V}}\right\|_{F}=$ 43.64 1.50 1.82 17.57 1.86
$\left\|{\bm{U}}^{\top}\Delta{\bm{W}}{\bm{V}}\right\|_{F}=$ 68.18 55.79 44.95
Amplification factor - - 37.46 3.18 24.17
Table 7: The Frobenius norm of ${\bm{U}}^{\top}{\bm{W}}{\bm{V}}$ where ${\bm{U}}$ and ${\bm{V}}$ are the left/right top $r$ singular vectors of either ${\bm{W}}$, a random matrix, or $\Delta{\bm{W}}$ of PEFTs. The weight matrices are taken from the query projection of the 16th layer of LLaMA2-7B.
Figure 2: Left: The subspace similarity averaged by all modules in a layer. Right: The subspace similarity averaged by all layers for a specific module type.
To further analyze the characteristics of $\Delta{\bm{W}}$ for different methods, we measure the normalized subspace similarity (Hu et al., 2021) between the subspaces spanned by top $r$ singular vectors of $\Delta{\bm{W}}$ and the subspace spanned by top $r$ singular vectors, bottom $r$ singular vectors, and a random $r$ singular vectors of ${\bm{W}}$. We define the normalized similarity metric based on the Grassmann distance following Hu et al. (2021) (See Appendix A.1 for more details). The results are presented in Figure 2. From the left figure, it is evident that PiSSA focuses on optimizing features associated with the top singular vectors, whereas LoRA and MiLoRA do not significantly optimize these features. MiLoRA primarily optimizes features related to the bottom singular vectors. The right figures further show that MiLoRA places greater emphasis on features associated with the bottom singular vectors of the MLP layers.
5.4 MiLoRA Forgets Less than Baselines
Method LoRA PiSSA MiLoRA
Forgetting loss 3.24 6.07 2.54
Table 8: Forgetting loss of various PEFT methods.
A hypothesis explanation for why MiLoRA outperforms baselines is that pre-trained knowledge is more retained. To test this hypothesis, we follow Kalajdzievski (2024) and use cross-entropy as the metric for measuring forgetting. This is the usual next token prediction loss used when training LLMs, except that the target next token is replaced by the distribution predicted by the pre-trained base model. We evaluate the forgetting metric on the LLaMA2-7B model finetuned in math reasoning using the WikiText-103 test dataset following Kalajdzievski (2024). As shown in Table 8, MiLoRA exhibits the lowest forgetting loss, which is consistent with our hypothesis that MiLoRA makes the least modification to pre-trained knowledge. In contrast, the loss of PiSSA is significantly higher than that of the other methods.
5.5 Comparison Between MiLoRA and PiSSA
Hyp. Setup PEFT GSM8K MATH Avg.
PiSSA LoRA 41.5 5.8 23.6
PiSSA 51.3 7.6 29.4
MiLoRA 40.0 5.2 22.6
LLM-Adapters LoRA 56.6 10.8 33.7
PiSSA 51.3 10.4 30.8
MiLoRA 58.6 11.6 35.1
Table 9: Math reasoning evaluation results for LLaMA2-7B with PiSSA hyperparameters and our (LLM-Adapters) hyperparameters.
Concurrent with our research, Meng et al. (2024) have recently introduced a low-rank adaptation method called PiSSA. PiSSA shares a similar framework as MiLoRA, but adapts the principal singular components. We argue that PiSSA and MiLoRA are fundamentally different.
Motivation: PiSSA is designed to approximate full finetuning by adapting the principal singular components, which are believed to capture the essence of the weight matrices. In contrast, our method MiLoRA aims to adapt to new tasks while maximally retaining the base model’s knowledge. To achieve this, we instead finetune the minor singular components of the weight matrices, which are less important for the weight matrices.
Performance: The PiSSA paper claims that PiSSA outperforms both LoRA and adapting minor singular components. However, our investigation suggests that this claim is likely due to the specific hyperparameters used in their experiments. To verify this, we replicate their experimental setup using the same MetaMathQA 100K dataset and compare the performance of PiSSA, LoRA, and MiLoRA under both PiSSA’s hyperparameters and our own. Specifically, we choose a rank of 128 where PiSSA obtains the best performance in their paper and finetunes for 1 epoch for all setups following PiSSA paper. The results are summarized in Table 9. We have several observations. First, under the PiSSA hyperparameter setup, PiSSA does indeed outperform both LoRA and MiLoRA, consistent with the claims made in the original PiSSA paper. Second, we find that all methods perform better in our hyperparameter setup, suggesting that our hyperparameter configuration is more effective. PiSSA configuration of hyperparameters has a combination of a small learning rate (2e-5) and a large batch size (128). This could potentially result in slow learning speed and sub-optimal performance for common PEFTs, as the optimal learning rate for common PEFTs is generally much higher than that for full finetuning (Lialin et al., 2023; Biderman et al., 2024). Third, in our hyperparameter setup, MiLoRA performs the best, outperforming LoRA and PiSSA by 1.4 and 4.3 average scores, respectively. In summary, when using an appropriate hyperparameter configuration, our proposed MiLoRA method achieves markedly superior performance compared to PiSSA.
6 Related Work
Parameter-Efficient Finetuning of LLMs.
After pretraining, LLMs are finetuned with instruction data or human preference data to adapt to different downstream tasks (Ouyang et al., 2022; Yu et al., 2023; Cui et al., 2023). Parameter-efficient finetuning (Lialin et al., 2023, PEFT) explores effective approaches to reduce the computational resources required during finetuning while managing to match the performance of full finetuning. Previous PEFT methods can be grouped into three lines: adapter-, LoRA- and prompt-based methods. Adapter-based methods insert additional trainable feedforward submodules into each Transformer layer (Houlsby et al., 2019; Pfeiffer et al., 2021a, b). However, they introduce additional computational costs during inference. LoRA-based methods (Hu et al., 2021; Liu et al., 2024b; Zhang et al., 2023) model the changes of selected linear weights as low-rank matrices. During finetuning, given a linear matrix, only two low-rank matrices are optimized while the pretrained weight matrix is frozen. Prompt-based methods (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2022) add additional soft prompts to the input tokens. During training, only soft prompts are updated while the pretrained model parameters are fixed. Among these PEFT methods, the LoRA-based approaches are widely used for LLM fine-tuning because they are easy to implement and do not introduce computational overhead during inference.
LoRA and Its Variants
LoRA (Hu et al., 2021) reparameterizes the weight update with two trainable low-rank matrices, while freezing the pretrained weights. With this lightweight decomposition, LoRA reduces storage and task-switching overhead by sharing the pretrained models across multiple tasks. Since then, researchers have explored and proposed different LoRA variants for PEFT. AdaLoRA (Zhang et al., 2023) and ALoRA (Liu et al., 2024d), adaptively determine the rank of LoRA module in each weight matrix according to the importance score. The rsLoRA (Kalajdzievski, 2023) modifies the LoRA with the appropriate scaling factor to improve the performance of large ranks. The DoRA (Liu et al., 2024b) method decomposes the pretrained weight into the magnitude and directional components, then finetunes both for better performance. Concurrent with our work, PISSA (Meng et al., 2024) and LoRA-GA Wang et al. (2024b) propose to better approximate full finetuning by only updating the principal singular components or aligning the gradients of low-rank updates to that of full finetuning.
7 Conclusion
In this paper, we introduce MiLoRA, a simple yet effective low-rank adaption method for LLMs. MiLoRA effectively learns on finetuning tasks while better preserving the pretrained knowledge by adapting the minor singular components of pretrained weight matrices. We investigate the effectiveness of MiLoRA on a wide range of LLM evaluation benchmarks, including commonsense reasoning, math reasoning and instruction-following, and vision-language model evaluation benchmarks. Experiment results demonstrate that MiLoRA consistently outperforms LoRA and PiSSA without sacrificing training or inference efficiency. We hope that our work will inspire future research on parameter-efficient finetuning of LLMs.
Limitations
Due to limited computational resources, we only examine the effectiveness of MiLoRA on the LLaMA-family LLMs and tasks like commonsense reasoning, math reasoning and instruction following benchmarks following previous work Liu et al. (2024b); Wu et al. (2024b). We examine and compare MiLoRA with baselines using the hyperparameter configurations provided in the previous works Hu et al. (2023); Liu et al. (2024b) instead of exhaustive hyperparameter search for each task. We leave the exploration of MiLoRA on other tasks and other LLMs like Mistral and Gemma as future work.
Appendix A Detailed Experiment Setups
A.1 Similarity Metric Between Subspaces
Following Hu et al. (2021), we use the measure $\phi({\bm{A}},{\bm{B}})=\frac{\left\|{\bm{A}}^{\top}{\bm{B}}\right\|_{F}^{2}}{r}$ to measure the similarity between two column orthonormal matrices ${\bm{A}},{\bm{B}}\in\mathbb{R}^{d\times r}$, where $F$ donotes the Frobenius norm. The value of $\phi({\bm{A}},{\bm{B}})$ ranges from 0 to 1, where 1 indicates complete overlap and 0 indicates complete separation.
Given two matrices $\Delta{\bm{W}}$ and ${\bm{W}}$, we extract $r$ left singular vectors from each to form the subspace matrices for them, which are denoted as $\Delta{\bm{W}}_{s}$ and ${\bm{W}}_{s}$ , and then the subspace similarity is computed using $\phi(\Delta{\bm{W}}_{s},{\bm{W}}_{s})$.
A.2 Our Hyperparameter Setup for LLM
Table 10 shows our detailed hyperparameters. This setup follows LLM-adapters (Hu et al., 2023), therefore we denote it as LLM-adapters setup.
Hyperparameters ComR MathR InsF
Rank r 32 64 32
$\alpha$ of LoRA 64 128 64
$\alpha$ of PiSSA/MiLoRA 32 64 32
Dropout 0.05
Optimizer AdamW
LR 3e-4
LR Scheduler Linear
Batch size 16
Warmup Steps 100
Epochs 3
Placement query, key, value, MLP up, MLP down
Table 10: Our hyperparameter configuration on the commonsense reasoning (ComR), math reasoning (MathR) and instruction-following (InsF) tasks.
Visual instruction tuning hyperparameters
Rank r 128
$\alpha$ of LoRA 256
$\alpha$ of PiSSA/MiLoRA 128
Dropout 0.05
Optimizer AdamW
LR 2e-4
LR Scheduler cosine
Batch Size 16
Warmup Ratio 0.03
Epochs 1
Placement query, key, value, output, gate, MLP up, MLP down
Table 11: Our hyperparameter configurations applied in visual instruction tuning tasks.
A.3 PiSSA Hyperparameter Setup for LLM
Table 12 shows the detailed hyperparameters from the PiSSA paper, which we denoted as PiSSA hyperparameter setup.
PiSSA hyperparameters
$\alpha$ Same as rank r
Dropout 0.0
Optimizer AdamW
LR 2e-5
LR Scheduler cosine
Batch Size 128
Warmup Ratio 0.03
Epochs 1
Placement query, key, value, output, gate, MLP up, MLP down
Table 12: The detailed hyperparameter configurations used by PiSSAMeng et al. (2024).
A.4 More Implementation Details
We use the implementation of LLM-Adapters (Hu et al., 2023) [^2] https://github.com/AGI-Edgerunners/LLM-Adapters for commonsense reasoning, the implementation of PiSSA (Meng et al., 2024) [^3] https://github.com/GraphPKU/PiSSA for math reasoning and the implementation of open-instruct [^4] https://github.com/allenai/open-instruct for instruction tuning. For hyperparameters of finetuning the vision-language model, we follow the LoRA configuration provided by Liu et al. (2024a) without tuning for all PEFT methods for a fair comparison. Table 11 shows the details.
Appendix B More Experiment Results
$\left\|{\bm{W}}\right\|_{F}=123.0$ ${\bm{W}}$ Random LoRA PiSSA MiLoRA
$\left\|{\bm{U}}^{\top}{\bm{W}}{\bm{V}}\right\|_{F}=$ 33.34 1.17 1.32 6.98 1.29
$\left\|{\bm{U}}^{\top}\Delta{\bm{W}}{\bm{V}}\right\|_{F}=$ 77.02 74.34 56.61
Amplification factor - - 58.35 10.65 43.88
Table 13: The Frobenius norm of ${\bm{U}}^{\top}{\bm{W}}{\bm{V}}$ where ${\bm{U}}$ and ${\bm{V}}$ are the left/right top $r$ singular vectors of either ${\bm{W}}$, a random matrix, or $\Delta{\bm{W}}$ of PEFTs. The weight matrices are taken from the MLP down projection of the 16th layer of LLaMA2-7B.
We present additional results about the relationship between $\Delta{\bm{W}}$ and ${\bm{W}}$ in Table 13, in which the weight matrices are taken from the MLP down projection of the 16th layer of LLaMA2-7B.
Appendix C Broader Impacts
The MiLoRA method enhances model performance with lower training costs. Teams or users with limited computational resources can finetune large models using MiLoRA, promoting the broader application of large models across diverse groups. However, the PEFT methods could potentially be exploited to finetune models for malicious purposes. It is essential to investigate rules and strategies to prevent the misuse of PEFT methods.